# Heath–Jarrow–Morton meet lifted Heston in energy markets for joint historical and implied calibration

### v1

@@ Abstract
In energy markets, joint historical and implied calibration is of paramount importance for practitioners yet notoriously challenging due to the need to align historical correlations of futures contracts with implied volatility smiles from the option market. We address this crucial problem with a parsimonious multiplicative multi-factor Heath-Jarrow-Morton (HJM) model for forward curves, combined with a stochastic volatility factor coming from the Lifted Heston model. We develop a sequential fast calibration procedure leveraging the Kemna-Vorst approximation of futures contracts: (i) historical correlations and the Variance Swap (VS) volatility term structure are captured through Level, Slope, and Curvature factors, (ii) the VS volatility term structure can then be corrected for a perfect match via a fixed-point algorithm, (iii) implied volatility smiles are calibrated using Fourier-based techniques. Our model displays remarkable joint historical and implied calibration fits - to both German power and TTF gas markets - and enables realistic interpolation within the implied volatility hypercube.
@@ Introduction
Several approaches to energy market modeling are considered in the literature, differing in the choice of the initial stochastic quantity to be modeled. The first class of models focuses on the spot price process, as seen in works like Mishura, Ottaviano, and Vargiolu (2023); Schmeck and Schwerin (2021); Cortazar, Lopez, and Naranjo (2017). Another approach, inspired by the LIBOR market model , see Brigo and Mercurio (2006), models futures contracts with specific delivery periods (e.g., monthly contracts, as in Kiesel, Schindlmayr, and Börger (2009), Gardini and Santilli (2024)), using these contracts as building blocks to derive the prices of other contracts under no-arbitrage conditions. The third class of models, inspired by the well-known Heath, Jarrow, and Morton (1992) (HJM) interest rate model, takes infinitesimal futures contracts as a starting point and uses them to reconstruct futures contracts with any delivery period. This type of model ensures consistent dynamics across all futures contracts while also guaranteeing the absence of arbitrage between futures contracts with overlapping delivery periods.
@@
The most significant limitation of the models mentioned above is that they allow for either historical calibration or implied calibration, but not both simultaneously. Hence the following questions:
@@
If so, can we do it with a parsimonious and tractable model?
We answer both questions affirmatively with a parsimonious multiplicative multi-factor Heath, Jarrow, and Morton (1992) (HJM) model for forward curves, combined with a stochastic volatility factor coming from the Lifted Heston model of Abi Jaber (2019).
@@
It is worth noting that these three steps are decoupled, done sequentially and made highly tractable thanks to the KV approximation. Our model displays remarkable joint historical and implied calibration fits to both German power and TTF gas markets. In order to validate our calibration methodology, we show a posteriori how close the KV approximated futures are to the exact arbitrage-free futures in terms of sample trajectories, implied volatility smiles and correlations between futures contracts. Finally, we show that such a fully calibrated model can be used to interpolate the implied volatility hypercube in a consistent manner. The main calibration results are collected in Section 5. Additional model algorithmic insights and calibration results are postponed to the appendices.
@@
For $N\in\mathbb{N}^{*}$, we denote by $\mathbb{S}_{++}^{N}$ (resp. $\mathbb{S}_{+}^{N}$) the set of $N\times N$ definite (resp. semi-definite) positive matrices, and .$\|.\|_{w}$ denotes a weighted Euclidean norm with weights $w:=\left(w_{i}\right)_{i\in\{1,\ldots,N\}^{2}}\in\mathbb{R}_{+}^{N}$ such that $\|u\|_{w}:=\sqrt{\sum_{i=1}^{N}w_{i}u_{i}^{2}},\;u\in\mathbb{R}^{N}$. We define similarly the weighted Frobenius norm $\|A\|_{\Gamma}:=\sqrt{\sum_{i=1}^{N}\sum_{j=i}^{N}\Gamma_{i,j}A_{i,j}^{2}},\;A\in\mathbb{R}^{N\times N}$, with matrix weights $\Gamma:=\left(\Gamma_{i,j}\right)_{i,j\in\{1,\ldots,N\}^{2}}\in\mathbb{R}_{+}^{N\times N}$. We will omit the indices $w$ and $\Gamma$ for standard Euclidean and Frobenius norms $\|u\|$ and $\|A\|$ respectively. We will also denote by $\|\cdot\|_{\infty}$ the infinity norm $\|u\|_{\infty}:=\max_{i\in\{1,\ldots,N\}}|u_{i}|$ and $\|A\|_{\infty}:=\max_{i,j\in\{1,\ldots,N\}}|A_{ij}|$.
@@ 1 The model: HJM with lifted Heston
  • $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V$ is a stochastic variance process responsible for the implied smile of the form
@@
    where the factors $(U^{i})_{i\in\{1,\ldots,M\}}$, weighted by $c_{i}\geq 0$, are driven by the same Brownian motion $B$, but mean-revert at different speeds $0<x_{1}<x_{2}<\ldots<x_{M}$ such that
@@
We prove in the next theorem that the model (1.1) is indeed well-defined leveraging results from Abi Jaber (2019). In particular, we note that although the different factors $U^{i}$ can become negative, the variance process $V$ is always nonnegative, as illustrated on Figure 1. Furthermore, the variance process (1.2) is Markovian in the state variables $U:=(U^{i})_{i\in\{1,\ldots,M\}}$ with a state space corresponding to the set of $u\in\mathbb{R}^{M}$ such that
@@
[…]
@@
$\dfrac{{\rm d}f(t,T)}{f(t,T)}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(t)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{V_{t}}}\sqrt{\sigma(t,T)^{\top}R\sigma(t,T)}{\rm d}\tilde{W}_{t},\quad d\langle B,\tilde{W}\rangle_{t}=\sigma(t,T)^{\top}\tilde{\rho}\,{\rm d}t,\quad t\in[0,T],$ (1.8)
@@
In practice, the infinitesimal futures contracts given by (1.1) are not observed in the market and therefore must be related to market traded futures contracts. Using practitioners’ vocabulary, “absolute” futures contracts delivering electricity on a fixed calendar delivery period $[T_{s},T_{e}]$ are typically quoted in organized markets until a few days before their first delivery date i.e. for dates $t$ such that $0\leq t\leq T_{s}-\delta$, $\delta$ equal to a few days. We recall in Section 2.1 the typical futures quoting in power markets. From the model perspective, the unitary absolute futures contract delivering continuously a unitary power unit of electricity over $[T_{s},T_{e}]$ is given by [^1] We assume a zero discount rate ($r=0$) for simplicity.
@@
1.1 The Kemna-Vorst approximation
@@
From now on, we fix a delivery period $[T_{s},\,T_{e}]$ and use the approximation $\widetilde{F}$ everywhere instead of the exact dynamics. Accordingly, we also use the notation $\Sigma_{.}:=\Sigma_{.}(T_{s},T_{e})$, leaving the delivery period implicit when there is no ambiguity.
@@ 1.2 Variance swap price and volatility
The variance swap volatility is a zero-order approximation of the ATM volatility level as shown by Bergomi and Guyon (2011), allowing us to formalize the idea of the implied volatility smile level.
@@ 1.3 A Nelson-Siegel parametrization for the volatility functions $\sigma$
as an alternative to the previously used polynomial fitting techniques to match the yields of US Treasury bills of maturity $m$, obtained in their formulation by integrating from zero to $m$ the forward rate (1.23) and dividing by $m$. They showed that such expression of the forward rate has a desirable non-explosive asymptotic behavior, and is capable of reproducing humps, $S$-shapes, and monotonic curves.
@@ 1.4 Joint calibration: overview and snapshots
Step 1: Joint historical and implied VS term structure calibrationData: Historical covariances and variance swapsStep 2: VS term structure fit correctionCalibrated $(\sigma,R)$Data: Variance swapsStep 3: Implied smile calibrationCalibrated $(g,h)$Data: Call and put optionsOutput: Calibrated $(\sigma,R,g,h,c,x,\tilde{\rho})$Calibrated $(c,x,\tilde{\rho})$
@@
In Figure 4, we provide a calibration snapshot that summarizes the results of all three calibration steps for the case of the German power market as of the $1^{\text{st}}$ of July 2024. For this date, $N=5$ $L$-$S$-$C$ risk factors with $N_{s}=3$, $N_{c}=1$, and $M=3$ stochastic volatility factors were used, which yields 19191919222Our choice of 19 parameters (5 factors) remains conservative compared to the literature on historical calibration in power markets: Féron and Gruet (2024) consider 5 $S$-factors (20 parameters), Gardini and Santilli (2024) use 10 factors with $100$ parameters to fit the realized volatility of $144$ futures from six different markets, and a PCA study in Koekebakker and Ollmar (2005) supports a similar order of factors. (resp. $11$) calibrated parameters for step 1 (resp. step 3). Our model achieves an excellent fit of the historical volatilities and correlations in step 1, a perfect fit of the VS term structures in step 2 and a remarkable fit of the whole implied volatility surface in step 3, see also Figure 15. More detailed specification of model factors and calibrated parameters will be described further in Section 5.1.

### v2

# 11affiliationtext: École Polytechnique, CMAP22affiliationtext: Engie Global Markets33affiliationtext: Université Paris Cité, LPSM

@@ Abstract
In energy markets, joint historical and implied calibration is of paramount importance for practitioners, yet notoriously challenging due to the need to align historical correlations of futures contracts with implied volatility smiles from the option market. We address this crucial problem with a multiplicative multi-factor Heath-Jarrow-Morton (HJM) model for forward curves, combined with a stochastic volatility factor coming from the lifted Heston model. We develop a sequential fast calibration procedure leveraging the Kemna-Vorst approximation of futures contracts: (i) historical correlations and the Variance Swap (VS) volatility term structure are captured through Level, Slope, and Curvature factors, (ii) the VS volatility term structure can then be corrected for a perfect match via a fixed-point algorithm, (iii) implied volatility smiles are calibrated using Fourier-based techniques. The main advantage of the proposed calibration framework is the decoupling of the calibration steps: each step tackles a simpler calibration subproblem and guaranties that the previously optimized parameters remain unchanged. Our model displays remarkable joint historical and implied calibration fits on the German power market and enables realistic interpolation within the implied volatility hypercube.
@@ Introduction
Several approaches to energy market modeling are considered in the literature, differing in the choice of the initial stochastic quantity to be modeled. The first class of models focuses on the spot price process, as seen in works like Mishura, Ottaviano, and Vargiolu (2023); Schmeck and Schwerin (2021); Cortazar, Lopez, and Naranjo (2017). Another approach, inspired by the LIBOR market model , see Brigo and Mercurio (2006), models futures contracts with specific delivery periods (e.g., monthly contracts, as in Kiesel, Schindlmayr, and Börger (2009), Gardini and Santilli (2024)), using these contracts as building blocks to derive the prices of other contracts under no-arbitrage conditions. The third class of models, inspired by the well-known Heath, Jarrow, and Morton (1992) (HJM) interest rate model, takes infinitesimal futures contracts as a starting point and uses them to reconstruct futures contracts with any delivery period. This type of model, adapted to the energy-markets setting by Clewlow and Strickland (1999), ensures consistent dynamics across all futures contracts while also guaranteeing the absence of arbitrage between futures contracts with overlapping delivery periods. The ability to model the dynamics of the entire futures curve, that is, the joint evolution of all futures contracts, is essential for pricing and hedging derivatives that are sensitive to cross-contract correlations, such as swing and spread options (Carmona and Durrleman, 2003).
Since the HJM approach naturally leads to modeling the entire futures curve, it is common to work with stochastic processes taking values in infinite-dimensional Hilbert spaces. In particular, several classes of infinite-dimensional stochastic volatility models have been developed; see, for instance, Cox, Karbach, and Khedher (2022). Benth and Simonsen (2018) consider an infinite-dimensional Heston-type model, Benth and Paraschiv (2018) employ an NIG Lévy process as the stochastic driver, and Benth and Harang (2021) analyze infinite-dimensional Volterra-type dynamics. For a more systematic treatment of infinite-dimensional modeling, we refer the reader to the monograph by Benth and Krühner (2023).
@@
In energy markets, the limited liquidity of options and the relatively small set of underlyings for which they are traded make it impossible to calibrate pricing models using option prices alone, a marked contrast to more developed derivatives markets, such as equities or interest rates. Consequently, a crucial requirement for any pricing model in this low-liquidity environment is the ability to produce reasonable interpolations and extrapolations of option prices, not only across strikes and maturities but also across different futures contracts, in order to fill the missing regions of the implied volatility hypercube. Moreover, as discussed above, calibrating the correlation structure is essential, yet cannot be inferred from option prices. Historical calibration therefore becomes indispensable for practical implementation. At the same time, given the gradually increasing liquidity of options, the options market can no longer be neglected, both because of regulatory requirements and due to the arbitrage opportunities that arise under purely historical models. The most significant limitation of the approaches mentioned above is that they allow for either historical calibration or implied calibration, but not both simultaneously. Hence the following questions:
@@
If so, can we do it with a tractable model that provides a consistent interpolation of option prices?
We answer both questions affirmatively with a multiplicative multi-factor Heath, Jarrow, and Morton (1992) (HJM) model for forward curves, combined with a stochastic volatility factor coming from the lifted Heston model of Abi Jaber (2019). To the best of our knowledge, this paper is the first to address both calibration problems with a single model, taking into account the entire implied volatility smile.
@@
We emphasize that the calibration procedure is fully decoupled: the three steps are performed sequentially, with each step preserving the calibration results achieved in the previous ones. Moreover, thanks to both the KV approximation and the decoupled structure, each step remains highly tractable. Decomposing the calibration into simpler subproblems yields a procedure that is a priori faster and more efficient than calibrating all parameters simultaneously. Our model displays remarkable joint historical and implied calibration fits to both German power and TTF gas markets. In order to validate our calibration methodology, we show a posteriori how close the KV approximated futures are to the exact arbitrage-free futures in terms of sample trajectories, implied volatility smiles and correlations between futures contracts. Finally, we show that such a fully calibrated model can be used to interpolate the implied volatility hypercube in a consistent manner. The main calibration results are collected in Section 5. Additional model algorithmic insights and calibration results are postponed to the appendices.
@@
For $N\in\mathbb{N}^{*}$, we denote by $\mathbb{S}_{++}^{N}$ (resp. $\mathbb{S}_{+}^{N}$) the set of $N\times N$ definite (resp. semi-definite) positive matrices, and $\|.\|_{w}$ denotes a weighted Euclidean norm with weights $w:=\left(w_{i}\right)_{i\in\{1,\ldots,N\}^{2}}\in\mathbb{R}_{+}^{N}$ such that $\|u\|_{w}:=\sqrt{\sum_{i=1}^{N}w_{i}u_{i}^{2}},\;u\in\mathbb{R}^{N}$. We define similarly the weighted Frobenius norm $\|A\|_{\Gamma}:=\sqrt{\sum_{i=1}^{N}\sum_{j=i}^{N}\Gamma_{i,j}A_{i,j}^{2}},\;A\in\mathbb{R}^{N\times N}$, with matrix weights $\Gamma:=\left(\Gamma_{i,j}\right)_{i,j\in\{1,\ldots,N\}^{2}}\in\mathbb{R}_{+}^{N\times N}$. We will omit the indices $w$ and $\Gamma$ for standard Euclidean and Frobenius norms $\|u\|$ and $\|A\|$ respectively. We will also denote by $\|\cdot\|_{\infty}$ the infinity norm $\|u\|_{\infty}:=\max_{i\in\{1,\ldots,N\}}|u_{i}|$ and $\|A\|_{\infty}:=\max_{i,j\in\{1,\ldots,N\}}|A_{ij}|$.
@@ 1 The model: HJM with lifted Heston
[…]
@@
    where the pseudo-factors $(U^{i})_{i\in\{1,\ldots,M\}}$, weighted by $c_{i}\geq 0$, are driven by the same Brownian motion $B$, but mean-revert at different speeds $0<x_{1}<x_{2}<\ldots<x_{M}$ such that
@@
We prove in the next theorem that the model (1.1) is indeed well-defined leveraging results from Abi Jaber (2019). In particular, we note that although the different factors $U^{i}$ can become negative, the variance process $V$ is always nonnegative (Abi Jaber et al., 2019, Theorem 7.1), as illustrated on Figure 1. Furthermore, the variance process (1.2) is Markovian in the state variables $U:=(U^{i})_{i\in\{1,\ldots,M\}}$ with a state space corresponding to the set of $u\in\mathbb{R}^{M}$ such that
@@
Remark 1.1.
We chose the lifted Heston model for two reasons. First, its affine structure allows for fast vanilla option pricing using Fourier techniques, which is highly beneficial for calibration speed. Second, it offers significantly more flexibility in the shapes of implied volatility surfaces that can be calibrated, compared to the simpler standard Heston model, as we demonstrate in Appendix A.1. Thus, the lifted Heston model strikes a good balance between numerical tractability and universality. However, we emphasize that the modeling framework we propose is not limited to this model and is flexible enough to accommodate any stochastic variance process $V_{t}$ satisfying ${\mathbb{E}}[V_{t}]\equiv 1$.
Figure 1: Trajectory of the stochastic variance factors $(U^{i})_{i=1,2,3}$ (on the left) and the resulting stochastic variance $V$ (on the right) corresponding to the parameters $x=(4.6\cdot 10^{-6},\,9.712,\,20.249)$ and c=(0.492, 0.68, 2.79)c=(0.492,\,0.68,\,2.79). The simulation scheme is described in Appendix D.
@@
$\dfrac{{\rm d}f(t,T)}{f(t,T)}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(t)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{V_{t}}}\sqrt{\sigma(t,T)^{\top}R\sigma(t,T)}{\rm d}\tilde{W}_{t},\quad d\langle B,\tilde{W}\rangle_{t}=\dfrac{\sigma(t,T)^{\top}\tilde{\rho}\,{\rm d}t}{{\sqrt{\sigma(t,T)^{\top}R\sigma(t,T)}}},\quad t\in[0,T],$ (1.8)
@@
In practice, the infinitesimal futures contracts given by (1.1) are not observed in the market and therefore must be related to market traded futures contracts. Using practitioners’ vocabulary, “absolute” futures contracts delivering electricity on a fixed calendar delivery period $T_{s},T_{e}]$ are typically quoted in organized markets until a few days before their first delivery date i.e. for dates $t$ such that $0\leq t\leq T_{s}-\delta$, $\delta$ equal to a few days. We recall in Section 2.1 the typical futures quoting in power markets. From the model perspective, the unitary absolute futures contract delivering continuously a unitary power unit of electricity over $T_{s},T_{e}]$ is given by [^1] We assume a zero discount rate ($r=0$) for simplicity.
@@
Remark 1.4.
In the discussion above, we assumed that the risk-free rate is $r=0$. In the general case, however, a similar no-arbitrage argument (Benth and Koekebakker, 2008, Section 4) yields
$F_{t}(T_{s},T_{e}):=\int_{T_{s}}^{T_{e}}w(T)\,f(t,T)\,\mathrm{d}T,\qquad 0\leq t<T_{s},$
where $w(T):=\frac{re^{-rT}}{e^{-rT_{s}}-e^{-rT_{e}}}.$ Thus, the extension of our approach to a non-zero constant interest rate is straightforward, although it makes the resulting expressions and computations more cumbersome. For notational convenience, we will therefore assume throughout the paper that $r=0$.
1.1 Correlation structure in the model
One of the main objectives of the HJM modeling framework is the consistent joint simulation of multiple futures contracts. This capability is essential for pricing derivatives that depend on several underlyings and are sensitive to their correlations, such as spread options. Consequently, the correlation structure between futures contracts plays a crucial role in both pricing and hedging such derivatives and should therefore be carefully calibrated.
We focus on the calibration of rolling covariances, that is, the covariances between returns of rolling futures contracts, defined as
$\mathrm{Cov}\left(r_{t}^{k_{1}}(\tau_{d}),\,r_{t}^{k_{2}}(\tau_{d})\right),\quad\text{where}\quad r_{t}^{k}(\tau_{d}):=\log\frac{{F}_{t}(t+T_{s}^{k},t+T_{e}^{k})}{{F}_{t-\tau_{d}}(t+T_{s}^{k},t+T_{e}^{k})},$ (1.14)
where the indices $k_{1}$ and $k_{2}$ correspond to the rolling contracts under consideration, and $\tau_{d}$ denotes the time step used for computing returns, typically set to one day.
The main motivation for calibrating rolling covariances is that they are assumed to be approximately constant over time and can therefore be estimated from historical data. The computation of rolling covariances in the model will be discussed in Section 3, while the historical estimation procedure is outlined in Appendix B.1.
However, although rolling covariances are convenient for calibrating the covariance structure of historical factors, pricing problems require dealing with instantaneous covariances of absolute contracts, given by
$\rho_{ij}(t)=\frac{\left\langle{\rm d}\log{F}(T_{s}^{i},T_{e}^{i}),\,{\rm d}\log{F}(T_{s}^{j},T_{e}^{j})\right\rangle_{t}}{\sqrt{\langle{\rm d}\log{F}(T_{s}^{i},T_{e}^{i})\rangle_{t}}\,\sqrt{\langle{\rm d}\log{F}(T_{s}^{j},T_{e}^{j})\rangle_{t}}},$ (1.15)
which have a direct impact on option prices, but cannot be estimated historically. The numerical precision of computing these correlations is discussed in Subsection 5.2.
We now turn to an approximation that renders the model tractable and, in particular, allows us to derive closed-form expressions for (1.14) and (1.15).
1.2 The Kemna–Vorst approximation
@@
Since the numerical analysis of the Kemna–Vorst approximation for our model, presented in Subsection 5.2, demonstrates that the approximated HJM model provides accurate pathwise and weak approximations and preserves the correlation structure well for contracts with the delivery periods of interest (i.e., one month, three months, and one year), we can use the approximated model interchangeably with the postulated model, thereby leveraging its tractability and analytical properties.
From now on, we fix a delivery period $T_{s},\,T_{e}]$ and use the approximation $\widetilde{F}$ everywhere instead of the exact dynamics. Accordingly, we also use the notation $\Sigma_{.}:=\Sigma_{.}(T_{s},T_{e})$, leaving the delivery period implicit when there is no ambiguity.
@@ 1.3 Variance swap price and volatility
For more profound discussion of variance swaps and their properties, we refer reader to Demeterfi et al. (1999).
@@
The variance swap volatility is a zero-order approximation of the ATM volatility level as shown by Bergomi and Guyon (2012), allowing us to formalize the idea of the implied volatility smile level. Moreover, formula (1.24) coincides exactly with the implied volatility for maturity $T$ in the HJM model under deterministic volatility, i.e., when $V\equiv 1$.
@@
We emphasize that our interest is not in variance swap contracts themselves, nor in hedging volatility risk; rather, we use the quantity (1.24), which naturally generalizes the implied volatility in the deterministic volatility case, only to calibrate the implied volatility term structure.
Indeed, under deterministic volatility, the quantity (1.23) reduces to the integrated variance up to maturity $T$, which was used, for instance, in Kiesel et al. (2009, Section 4) to calibrate the implied volatility term structure. This calibration procedure corresponds to the first step of our joint calibration. However, unlike Kiesel et al. (2009) and other classical HJM calibration approaches, in the second step we also calibrate the functions $g$ and $h$ to achieve an exact fit of the term structure, and in the third step we calibrate the implied volatility smile while keeping the variance swap volatility fixed.
@@ 1.4 A Nelson-Siegel parametrization for the volatility functions $\sigma$
as an alternative to the previously used polynomial fitting techniques to match the yields of US Treasury bills of maturity $m$, obtained in their formulation by integrating from zero to $m$ the forward rate (1.25) and dividing by $m$. They showed that such expression of the forward rate has a desirable non-explosive asymptotic behavior, and is capable of reproducing humps, $S$-shapes, and monotonic curves. We refer to Sepp and Rakhmonov (2023) as an example of a recent application of the HJM approach with Nelson–Siegel parametrization and stochastic volatility to the interest rate market.
@@ 1.5 Joint calibration: overview and snapshots
[…]
@@
In Figure 4, we provide a calibration snapshot that summarizes the results of all three calibration steps for the case of the German power market as of the $1^{\text{st}}$ of July 2024. For this date, $N=5$ $L$-$S$-$C$ risk factors with $N_{s}=3$, $N_{c}=1$, and $M=3$ stochastic volatility factors were used, which yields 1919222Our choice of 19 parameters (5 factors) remains conservative compared to the literature on historical calibration in power markets: Féron and Gruet (2024) consider 5 $S$-factors (20 parameters), Gardini and Santilli (2024) use 10 factors with $100$ parameters to fit the realized volatility of $144$ futures from six different markets, and a PCA study in Koekebakker and Ollmar (2005) supports a similar order of factors. (resp. $11$) calibrated parameters for step 1 (resp. step 3). Our model achieves an excellent fit of the historical volatilities and correlations in step 1, a perfect fit of the VS term structures in step 2 and a remarkable fit of the whole implied volatility surface in step 3, see also Figure 15. More detailed specification of model factors and calibrated parameters will be described further in Section 5.1.